Determine the uncorrected and corrected weights based on the following data:


A.Fine Aggregates - Natural

Fineness Modulus = 2.80

Specific Gravity (SSD) = 2.50

Moisture Content,% = 0.50

Absorption, % = 1.80


B.Coarse Aggregates - Angular

Maximum size,mm = 19.00

Spec.Gravity(SSD) = 2.65

Moisture Content,% = 1.35

Absorption,% = 0.80


C.Class of Concrete = Structural


D.Cement Factor = 11.10 bags/cu.m.


E.Slump,mm = 76.20


F.Spec. Gravity of Cement = 3.15


G.Flexural Strength,Mpa = 27.50


1
Expert's answer
2022-03-29T10:51:02-0400

corrected weight = aggregate weight "\times" 105%


= 0.50 "\times" 104.45%


= 52.225


uncorrected weight => "Wc = Wo\times(\sqrt1+ (\frac{s}{100})^2)"


"Wc = 76.20\times(\sqrt1+ (\frac{19}{100})^2)"


= 78.95